# DiCoRe: Enhancing Zero-shot Event Detection via Divergent-Convergent LLM Reasoning

### v1

@@ C.2 Complete Results for Transfer Learning Baselines
C.3 Complete Results for Reasoning Baselines
@@
Table 12: Complete results for comparison of DiCoRe with other fine-tuned transfer-learning approaches for zero-shot ED. *Training done for models other than DiCoRe. DiCoRe results are pure zero-shot, i.e., without any training. "-" indicates training data or where results were degenerate. (XX) = number of distinct event types.
@@
Table 14: Complete results for comparison of DiCoRe with reasoning approaches like Chain-of-thought (CoT) and thinking-based models for zero-shot ED. bold = best performance. (XX) = number of distinct event types.
@@
C.4 Additional results for Ablation Study

### v2

@@ C.4 Complete Results for Transfer Learning Baselines
C.5 Complete Results for Reasoning Baselines
@@
Table 14: Complete results for comparison of DiCoRe with other fine-tuned transfer-learning approaches for zero-shot ED. *Training done for models other than DiCoRe. DiCoRe results are pure zero-shot, i.e., without any training. "-" indicates training data or where results were degenerate. ($X$) = number of distinct event types.
@@
Table 16: Complete results for comparison of DiCoRe with reasoning approaches like Chain-of-thought (CoT) and thinking-based models for zero-shot ED. bold = best performance. ($X$) = number of distinct event types.
@@
C.6 Additional results for Ablation Study